OpenStudy (anonymous):

tickets for a concert cost $30 each. you have $210 to spend. write an axpression for the money that you have left after puchasing t tickets. Evaluate the expression to find the amount of money left after purchasing 6 tickets. answer 1___________ answer 2 ________________

OpenStudy (anonymous):

30t=x 30(6)=210

OpenStudy (anonymous):

there should be 2 answers

OpenStudy (anonymous):

1. Let x be the number of tickets, then the amount of money left, in dollars, is given by the function\[f(x)=210-30x\] 2. Replace x with 6\[f(x)=210-30\times 6=\]\[f(x)=210-180=\]\[30\]So, after buying 6 tickets you have 30$ left.